\begin{thebibliography}{10}

\bibitem{darwin1998expression}
Charles Darwin and Phillip Prodger.
\newblock {\em The expression of the emotions in man and animals}.
\newblock Oxford University Press, USA, 1998.

\bibitem{kollias2017recognition}
Dimitrios Kollias, Mihalis~A Nicolaou, Irene Kotsia, Guoying Zhao, and Stefanos
  Zafeiriou.
\newblock Recognition of affect in the wild using deep neural networks.
\newblock In {\em Proceedings of the IEEE Conference on Computer Vision and
  Pattern Recognition Workshops}, pages 26--33, 2017.

\bibitem{kollias2019deep}
Dimitrios Kollias, Panagiotis Tzirakis, Mihalis~A Nicolaou, Athanasios
  Papaioannou, Guoying Zhao, Bj{\"o}rn Schuller, Irene Kotsia, and Stefanos
  Zafeiriou.
\newblock Deep affect prediction in-the-wild: Aff-wild database and challenge,
  deep architectures, and beyond.
\newblock {\em International Journal of Computer Vision}, pages 1--23, 2019.

\bibitem{zafeiriou2017aff}
Stefanos Zafeiriou, Dimitrios Kollias, Mihalis~A Nicolaou, Athanasios
  Papaioannou, Guoying Zhao, and Irene Kotsia.
\newblock Aff-wild: Valence and arousal ‘in-the-wild’challenge.
\newblock In {\em Computer Vision and Pattern Recognition Workshops (CVPRW),
  2017 IEEE Conference on}, pages 1980--1987. IEEE, 2017.

\bibitem{kollias2022abaw}
Dimitrios Kollias.
\newblock Abaw: Learning from synthetic data \& multi-task learning challenges.
\newblock {\em arXiv preprint arXiv:2207.01138}, 2022.

\bibitem{kollias2021distribution}
Dimitrios Kollias, Viktoriia Sharmanska, and Stefanos Zafeiriou.
\newblock Distribution matching for heterogeneous multi-task learning: a
  large-scale face study.
\newblock {\em arXiv preprint arXiv:2105.03790}, 2021.

\bibitem{kollias2021analysing}
Dimitrios Kollias and Stefanos Zafeiriou.
\newblock Analysing affective behavior in the second abaw2 competition.
\newblock In {\em Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pages 3652--3660, 2021.

\bibitem{kollias2021affect}
Dimitrios Kollias and Stefanos Zafeiriou.
\newblock Affect analysis in-the-wild: Valence-arousal, expressions, action
  units and a unified framework.
\newblock {\em arXiv preprint arXiv:2103.15792}, 2021.

\bibitem{kollias2020analysing}
D~Kollias, A~Schulc, E~Hajiyev, and S~Zafeiriou.
\newblock Analysing affective behavior in the first abaw 2020 competition.
\newblock In {\em 2020 15th IEEE International Conference on Automatic Face and
  Gesture Recognition (FG 2020)(FG)}, pages 794--800.

\bibitem{kollias2019expression}
Dimitrios Kollias and Stefanos Zafeiriou.
\newblock Expression, affect, action unit recognition: Aff-wild2, multi-task
  learning and arcface.
\newblock {\em arXiv preprint arXiv:1910.04855}, 2019.

\bibitem{kollias2019face}
Dimitrios Kollias, Viktoriia Sharmanska, and Stefanos Zafeiriou.
\newblock Face behavior a la carte: Expressions, affect and action units in a
  single network.
\newblock {\em arXiv preprint arXiv:1910.11111}, 2019.

\bibitem{sak2014long}
Ha{\c{s}}im Sak, Andrew Senior, and Fran{\c{c}}oise Beaufays.
\newblock Long short-term memory based recurrent neural network architectures
  for large vocabulary speech recognition.
\newblock {\em arXiv preprint arXiv:1402.1128}, 2014.

\bibitem{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin.
\newblock Attention is all you need.
\newblock {\em Advances in neural information processing systems}, 30, 2017.

\bibitem{iandola2014densenet}
Forrest Iandola, Matt Moskewicz, Sergey Karayev, Ross Girshick, Trevor Darrell,
  and Kurt Keutzer.
\newblock Densenet: Implementing efficient convnet descriptor pyramids.
\newblock {\em arXiv preprint arXiv:1404.1869}, 2014.

\bibitem{duta2021improved}
Ionut~Cosmin Duta, Li~Liu, Fan Zhu, and Ling Shao.
\newblock Improved residual networks for image and video recognition.
\newblock In {\em 2020 25th International Conference on Pattern Recognition
  (ICPR)}, pages 9415--9422. IEEE, 2021.

\bibitem{howard2019searching}
Andrew Howard, Mark Sandler, Grace Chu, Liang-Chieh Chen, Bo~Chen, Mingxing
  Tan, Weijun Wang, Yukun Zhu, Ruoming Pang, Vijay Vasudevan, et~al.
\newblock Searching for mobilenetv3.
\newblock In {\em Proceedings of the IEEE/CVF international conference on
  computer vision}, pages 1314--1324, 2019.

\bibitem{eyben2015geneva}
Florian Eyben, Klaus~R Scherer, Bj{\"o}rn~W Schuller, Johan Sundberg, Elisabeth
  Andr{\'e}, Carlos Busso, Laurence~Y Devillers, Julien Epps, Petri Laukka,
  Shrikanth~S Narayanan, et~al.
\newblock The geneva minimalistic acoustic parameter set (gemaps) for voice
  research and affective computing.
\newblock {\em IEEE transactions on affective computing}, 7(2):190--202, 2015.

\bibitem{schuller2016interspeech}
Bj{\"o}rn Schuller, Stefan Steidl, Anton Batliner, Julia Hirschberg, Judee~K
  Burgoon, Alice Baird, Aaron Elkins, Yue Zhang, Eduardo Coutinho, Keelan
  Evanini, et~al.
\newblock The interspeech 2016 computational paralinguistics challenge:
  Deception, sincerity \& native language.
\newblock In {\em 17TH Annual Conference of the International Speech
  Communication Association (Interspeech 2016), Vols 1-5}, pages 2001--2005,
  2016.

\bibitem{hershey2017cnn}
Shawn Hershey, Sourish Chaudhuri, Daniel~PW Ellis, Jort~F Gemmeke, Aren Jansen,
  R~Channing Moore, Manoj Plakal, Devin Platt, Rif~A Saurous, Bryan Seybold,
  et~al.
\newblock Cnn architectures for large-scale audio classification.
\newblock In {\em 2017 ieee international conference on acoustics, speech and
  signal processing (icassp)}, pages 131--135. IEEE, 2017.

\bibitem{baevski2020wav2vec}
Alexei Baevski, Yuhao Zhou, Abdelrahman Mohamed, and Michael Auli.
\newblock wav2vec 2.0: A framework for self-supervised learning of speech
  representations.
\newblock {\em Advances in Neural Information Processing Systems},
  33:12449--12460, 2020.

\bibitem{desplanques2020ecapa}
Brecht Desplanques, Jenthe Thienpondt, and Kris Demuynck.
\newblock Ecapa-tdnn: Emphasized channel attention, propagation and aggregation
  in tdnn based speaker verification.
\newblock {\em arXiv preprint arXiv:2005.07143}, 2020.

\bibitem{hsu2021hubert}
Wei-Ning Hsu, Benjamin Bolte, Yao-Hung~Hubert Tsai, Kushal Lakhotia, Ruslan
  Salakhutdinov, and Abdelrahman Mohamed.
\newblock Hubert: Self-supervised speech representation learning by masked
  prediction of hidden units.
\newblock {\em IEEE/ACM Transactions on Audio, Speech, and Language
  Processing}, 29:3451--3460, 2021.

\bibitem{he2022masked}
Kaiming He, Xinlei Chen, Saining Xie, Yanghao Li, Piotr Doll{\'a}r, and Ross
  Girshick.
\newblock Masked autoencoders are scalable vision learners.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and
  Pattern Recognition}, pages 16000--16009, 2022.

\bibitem{jiang2020dfew}
Xingxun Jiang, Yuan Zong, Wenming Zheng, Chuangao Tang, Wanchuang Xia, Cheng
  Lu, and Jiateng Liu.
\newblock Dfew: A large-scale database for recognizing dynamic facial
  expressions in the wild.
\newblock In {\em Proceedings of the 28th ACM International Conference on
  Multimedia}, pages 2881--2889, 2020.

\bibitem{DBLP:conf/cvpr/Benitez-QuirozS16}
Carlos~Fabian Benitez{-}Quiroz, Ramprakash Srinivasan, and Aleix~M.
  Mart{\'{\i}}nez.
\newblock Emotionet: An accurate, real-time algorithm for the automatic
  annotation of a million facial expressions in the wild.
\newblock In {\em 2016 {IEEE} Conference on Computer Vision and Pattern
  Recognition, {CVPR} 2016, Las Vegas, NV, USA, June 27-30, 2016}, pages
  5562--5570. {IEEE} Computer Society, 2016.

\bibitem{DBLP:conf/cvpr/WangSHLGZGZ22}
Yan Wang, Yixuan Sun, Yiwen Huang, Zhongying Liu, Shuyong Gao, Wei Zhang,
  Weifeng Ge, and Wenqiang Zhang.
\newblock Ferv39k: {A} large-scale multi-scene dataset for facial expression
  recognition in videos.
\newblock In {\em {IEEE/CVF} Conference on Computer Vision and Pattern
  Recognition, {CVPR} 2022, New Orleans, LA, USA, June 18-24, 2022}, pages
  20890--20899. {IEEE}, 2022.

\bibitem{BarsoumICMI2016}
Emad Barsoum, Cha Zhang, Cristian Canton~Ferrer, and Zhengyou Zhang.
\newblock Training deep networks for facial expression recognition with
  crowd-sourced label distribution.
\newblock In {\em ACM International Conference on Multimodal Interaction
  (ICMI)}, 2016.

\bibitem{li2017reliable}
Shan Li, Weihong Deng, and JunPing Du.
\newblock Reliable crowdsourcing and deep locality-preserving learning for
  expression recognition in the wild.
\newblock In {\em 2017 IEEE Conference on Computer Vision and Pattern
  Recognition (CVPR)}, pages 2584--2593. IEEE, 2017.

\bibitem{li2019reliable}
Shan Li and Weihong Deng.
\newblock Reliable crowdsourcing and deep locality-preserving learning for
  unconstrained facial expression recognition.
\newblock {\em IEEE Transactions on Image Processing}, 28(1):356--370, 2019.

\bibitem{mollahosseini2017affectnet}
Ali Mollahosseini, Behzad Hasani, and Mohammad~H Mahoor.
\newblock Affectnet: A database for facial expression, valence, and arousal
  computing in the wild.
\newblock {\em IEEE Transactions on Affective Computing}, 10(1):18--31, 2017.

\bibitem{kingma2014adam}
Diederik~P Kingma and Jimmy Ba.
\newblock Adam: A method for stochastic optimization.
\newblock {\em arXiv preprint arXiv:1412.6980}, 2014.

\end{thebibliography}

\pdfoutput=1

\documentclass[10pt,twocolumn,letterpaper]{article}

\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{booktabs}
\usepackage{cite}
\usepackage[accsupp]{axessibility}  

\usepackage[pagebackref,breaklinks,colorlinks]{hyperref}

\usepackage[capitalize]{cleveref}
\crefname{section}{Sec.}{Secs.}
\Crefname{section}{Section}{Sections}
\Crefname{table}{Table}{Tables}
\crefname{table}{Tab.}{Tabs.}

\date{}

\begin{document}

\title{Multi-modal Expression Recognition with Ensemble Method}

\author{
Chuanhe Liu\textsuperscript{1, $\ast$, $\dagger$}, Xinjie Zhang\textsuperscript{2, $\ast$}, Xiaolong Liu\textsuperscript{1}, Tenggan Zhang\textsuperscript{2}\\ 
Liyu Meng\textsuperscript{1}, Yuchen Liu\textsuperscript{2}, Yuanyuan Deng\textsuperscript{1}, Wenqiang Jiang\textsuperscript{1} \\
 \textsuperscript{1} \small{Beijing Seek Truth Data Technology Co.,Ltd.} \\
 \textsuperscript{2} \small{School of Information, Renmin University of China} \\
}
\maketitle


\def\thefootnote{$\ast$}\footnotetext{These authors contributed equally to this work and should be considered co-first authors.}
\def\thefootnote{$\dagger$}\footnotetext{Corresponding Author.}

\maketitle

\begin{abstract}
   This paper presents our submission to the Expression Classification Challenge of the fifth Affective Behavior Analysis in-the-wild (ABAW) Competition. 
   In our method, multimodal feature combinations extracted by several different pre-trained models are applied to capture more effective emotional information. For these combinations of visual and audio modal features, we utilize two temporal encoders to explore the temporal contextual information in the data. In addition, we employ several ensemble strategies for different experimental settings to obtain the most accurate expression recognition results. 
   Our system achieves the average F1 Score of $0.45774$ on the validation set.
\end{abstract}

\section{Introduction}
Affective computing has a wide spectrum of application requirements in human-computer interaction, security, robotics manufacturing, automation, medical, and communications. Actively creating machines that can understand the feelings, emotions, and behaviors of humans would help them interact with humans in a more intimate way and serve effectively\cite{darwin1998expression}. Facial expressions are one of the most powerful, natural, and pervasive signals that humans use to communicate their emotional state and intentions. Machines can analyze human expressions leading to understanding human emotions. The majority of recent research in emotion recognition is based on deep learning, which requires a large quantity of labeled data. Nowadays, there are several datasets, such as Aff-wild\cite{kollias2017recognition, kollias2019deep, zafeiriou2017aff} and Aff-wild2\cite{kollias2022abaw,kollias2021distribution,kollias2021analysing,kollias2021affect,kollias2020analysing,kollias2019expression,kollias2019face,kollias2019deep,zafeiriou2017aff}, which provide us with large-scale data with high-quality labels, which are convenient for training neural networks and increasing the accuracy of expression recognition.

The information of facial expressions can be mainly obtained from the visual modality. Nevertheless, it is well known that audio modality also contains certain emotional information. The information of unimodal modality may be affected by various noises. To obtain a more complete emotional state, information from multiple modalities can be utilized. The multiple modalities' information can supplement and enhance the information to a degree, improving the recognition ability, generalization, and robustness of the model.

Our system for the expression recognition challenge contains several key components. First, the officially provided cropped aligned image data along with the labels are corresponded to, and the data containing labels but no images are complemented. Second, multiple pre-trained feature extractors are employed to extract visual and audio features. Then, we designed multimodal feature combinations and concatenated multiple features into multimodal feature representations. Two different types of temporal encoders, LSTM\cite{sak2014long} and Transformer\cite{vaswani2017attention}, are applied to extract contextual information from the multimodal features. Several techniques are also utilized for optimization. Finally, we adopted several ensemble strategies to ensemble the experimental results for different settings to raise the accuracy of recognition.

\section{Method}

For a given video $X$, it can be separated into tuo parts, the visual data $X^{vis}$ and the audio data $X^{aud}$. The visual data can be stated as a image frames sequence$\{F_{1}, F_{2}, ..., F_{n}\}$, and $n$ denotes the number of image frames in $X$. The goal of the Expression
Classification Challenge is to predict the sentiment label for each frame in the video. 


\subsection{Pre-processing}
Firstly, the officially provided video data is divided into multiple image frames. For each image frame, the face and facial landmarks are recognized by the face detector. The face part is cropped out according to the bounding box to facilitate the extraction of more accurate emotional information later. In order to be consistent with the official labels provided, we use the cropped and aligned face images provided by the competition for the actual processing.

We matched the labels with the images one by one, and found that some of the images corresponding to the labels did not exist in the cropped and aligned set given by the competition, probably because the face images of the corresponding frames were not detected in the video due to the lighting, angle, and other circumstances. For each of these non-existent images, we complement it by finding the nearest frame in the temporal dimension.

\subsection{Multimodal Feature Representation}


During feature extracting phase, we extract multiple type feature in visual and audio modality. 
The model used for extracting visual features including DenseNet-based\cite{iandola2014densenet} model, IResNet100-based\cite{duta2021improved} model, IResNet100-based\cite{iandola2014densenet} facial action unit (FAU) detection model and the MobileNet-based\cite{howard2019searching} model.
The model used for extracting audio features including eGeMAPS\cite{eyben2015geneva}, ComParE 2016\cite{schuller2016interspeech}, VGGish\cite{hershey2017cnn}, Wav2Vec 2.0\cite{baevski2020wav2vec}, ECAPA-TDNN\cite{desplanques2020ecapa} and HuBERT\cite{hsu2021hubert}.

\subsubsection{Visual Features} 
The first type of visual features is extracted by a pre-trained DenseNet model. Specifically, the DenseNet model is pre-trained on the FER+ and the AffectNet datasets. The dimension of the DenseNet-based visual features is 342. And this kind of feature is denoted as densenet.

The second type of visual feature is MAE-based model\cite{he2022masked}. MAE is a self-supervised model that trains label-free data by masking random patches from the input image and reconstructing the missing patches in the pixel space. We used a face dataset of scale 1.2 million, including DFEW\cite{jiang2020dfew}, Emotionet\cite{DBLP:conf/cvpr/Benitez-QuirozS16}, FERV39k\cite{DBLP:conf/cvpr/WangSHLGZGZ22} and so on, to pre-train the MAE encoder. We denotes this kind of feature as mae, and the dimension of it is 768.

The third type of visual feature is IResNet100-based model, and the dimension of IResNet100-based feature is $512$. 
A large-scale facial expression recognition data which consists of FER+\cite{BarsoumICMI2016}, RAF-DB\cite{li2017reliable}\cite{li2019reliable} and AffectNet\cite{mollahosseini2017affectnet} dataset is utilized to pretrain our facial expression recognition IResNet100-based model, which is denoted as ires100.
And a commercial authorized facial action unit detection dataset is used to pretrain the other IResNet100-based model, which is denoted as fau.

The fourth type of visual feature is MobileNet-based model, and the dimension of MobieNet-based feature is $512$.
We utilized AffectNet\cite{mollahosseini2017affectnet} to train a valence-arousal prediction model in the valence-arousal estimation task of AffectNet.



\subsubsection{Audio Features} 

The first type of audio feature is hand-craft features, which consists of eGeMAPS, ComParE 2016 and fbank. eGeMAPS and ComParE 2016 can be extracted using openSmile, and the dimension of these features are 23 and 130. The dimension of fbank is 80. For convenience, we denotes them as egemaps, compare and fbank.

The second type of audio feature is deep features, which consists of Wav2Vec 2.0, ECAPA-TDNN, VGGish and HuBERT. The dimension of Wav2Vec 2.0 feature is $1024$, the dimension of ECAPA-TDNN feature is $512$, the dimension of VGGish feature is $128$ and the dimension of HuBERT is $512$. We denotes them as wav2vec, ecapatdnn, vggish and hubert respectively.

\subsubsection{Multimodal Fusion}
Given the visual features $f^{v}$ and audio features $f^{a}$ corresponding to a frame, they are first concatenated and then fed into a fully-connected layer to produce the multimodal features $f^{m}$. It can be formulated as follows:

\begin{small}
\begin{equation}
    \centering
    f^{m} = W_{f}[f^{v};f^{a}] + b_{f}
\end{equation}
\end{small}
where $W_{f}$ and $b_{f}$ are learnable parameters. 

%


\subsection{Temporal Encoder}

Due to the limitation of GPU memory, we split the videos into segments at first. Given the segment length $l$ and stride $p$, a video with $n$ frames would be split into $[n/p]+1$ segments, where the $i$-th segment contains frames $\{F_{(i-1)*p+1}, ..., F_{(i-1)*p+l}\}$. 
With the multimodal features of the $i$-th segment $f^{m}_{i}$, we employ a temporal encoder to model the temporal context in the video. Specifically, two kinds of structures are utilized as the temporal encoder, including LSTM and Transformer Encoder.

\subsubsection{LSTM-based Temporal Encoder}
Long and short term memory networks (LSTM) are commonly applied to model sequential dependencies of time sequences. In a practical game, we use LSTM to model the temporal relationships in a sequence of frame images from a video. For the $i$-th video segment $s_{i}$, the multimodal features $f^{m}_{i}$ are directly fed into the LSTM. In addition, the last hidden states of the previous segment $s_{i-1}$ are also fed into the LSTM to encode the context between two adjacent segments. It can be formulated as follows:
\begin{small}
\begin{equation}
    \centering
    g_{i}, h_{i} = \text{LSTM}(f^{m}_{i}, h_{i-1})
\end{equation}
\end{small}
where $h_{i}$ denotes the hidden states at the end of $s_{i}$. $h_{0}$ is initialized to be zeros. To ensure that the last frame of $s_{i-1}$ and the first frame of segment $s_{i}$ are consecutive frames, there is no overlap between two adjacent segments when LSTM is used as the temporal encoder. In another word, the stride $p$ is the same as the segment length $l$.

\subsubsection{Transformer-Based Temporal Encoder}

We used a Transformer Encoder to model the temporal feature in the video segment, which can be formulated as:

\begin{small}
\begin{equation}
    \centering
    g_{i} = \text{TRMEncoder}(f^{m}_{i})
\end{equation}
\end{small}

Unlike LSTM, the transformer encoder just models the context in a single segment and ignores the dependencies of frames between segments. 




\begin{table*}[]
\begin{center}
  \caption{The performance of our method on the validation set.}
  \label{tab:overall_performence}
\begin{tabular}{c|c|c|c}
\hline
Model              & Visual Features               & Audio Features                 & F1      \\ \hline
Transformer        & mae,ires100                   & wav2vec,ecapatdnn,hubert       & 0.38362 \\
Transformer        & mae,ires100,densenet          & ecapatdnn,hubert               & 0.39112 \\
Transformer        & mae,ires100,fau,densenet      & ecapatdnn,hubert               & 0.3938  \\ 
lstm               & mae,ires100                   & wav2vec,ecapatdnn,hubert       & 0.37972 \\
lstm               & fau,ires100                   & ecapatdnn,hubert               & 0.38928 \\
lstm               & mae,ires100                   & wav2vec                        & 0.39385 \\
lstm               & densenet,mae,ires100          & wav2vec,ecapatdnn,hubert       & 0.40178 \\
lstm               & densenet,mae,ires100          & ecapatdnn,hubert               & 0.41410\\ \hline
\end{tabular}
\end{center}
\end{table*}

\subsection{Loss Function}
In the training phase, we utilize the RDrop loss which can be formulated as
\begin{small}
    \begin{equation}
    \begin{aligned}
        \centering
        L^{EXPR} = & \frac{1}{2} * (CE(\hat{y}_{1}, y) + CE(\hat{y}_{2}, y)) + \\
        & \alpha * \frac{1}{2} * (KL(\hat{y}_{1}, \hat{y}_{2}) + KL(\hat{y}_{2}, \hat{y}_{1}))
    \end{aligned}
    \end{equation}
    \label{eq:loss}
\end{small}
where $CE$ denotes cross entropy loss, $KL$ denotes Kullback-Leibler divergence loss. $y_{1}$ and $y_{2}$ denotes the first and the second inference prediction logits, $y$ denotes the label.



\section{Experiments}
\subsection{Dataset} 


Expression(Expr) Classification Challenge in fifth ABAW competition is based on Aff-Wild2, a large-scale dataset. Aff-Wild2 is consists of 548 videos and is annotated with 8 expressions (i.e. neutral, anger, disgust, fear, happiness, sadness, surprise and other). 
As for the feature extractors, we used some other dataset for pretraining, which consists of FER+\cite{BarsoumICMI2016}, RAF-DB\cite{li2017reliable, li2019reliable} and AffectNet\cite{mollahosseini2017affectnet}. 
In addition, an authorized commercial FAU dataset is also used for pretraining visual feature extractor, which consists of 7K images in 15 face action unit categories(AU1, AU2, AU4, AU5, AU6, AU7, AU9, AU10, AU11, AU12, AU15, AU17, AU20, AU24, and AU26).
As for the audio feature extractors, we used some different open-source models to extract features. Wav2Vec 2.0\cite{baevski2020wav2vec}, HuBERT\cite{hsu2021hubert} and ECAPA-TDNN\cite{desplanques2020ecapa} are the deep open-source model for extracting audio features.

\subsection{Evaluation Metric}
According to the competition regulations, we use the average F1 score across 8 categories, which can be formulated as 
\begin{small}
\begin{equation}
    \label{eq:metric}
    \centering
    p = \frac{\sum_{i}^{8}F1(\hat{y}_{i}, y_{i})}{8}
\end{equation}
\end{small}
where $F1$ denotes F1 score, $\hat{y}_{i}$ and $y_{i}$ denotes the $i$-th category of prediction and label respectively.

\subsection{Experiment Settings}

First declare that we used Adam\cite{kingma2014adam} optimizer to train models for 25 epochs.
As for the Transformer model, learning rate is 0.0001, the $\alpha$ in equation\ref{eq:loss} is $5$, the affine dimension is $1024$, the number of Transformer encoder layers is $4$, attention heads number is $4$, the dropout ratio in Transformer encoder layer is $0.3$, the sequence length of one segement is $128$, and the hidden size of head layers are \{512, 256\}.

\begin{table}[]
\begin{center}
  \caption{The performance of our method on the 5-fold cross-validation. The First 4 folds are from train set, and the last fold is the original validation set.}
  \label{tab:5fold}
\begin{tabular}{c|c}
\hline
         & F1       \\ \hline
Fold 1   & 0.43697  \\
Fold 2   & 0.38015  \\
Fold 3   & 0.35646  \\
Fold 4   & 0.39170  \\
Fold 5   & 0.38146  \\
Average  & 0.38935  \\ \hline
\end{tabular}
\end{center}
\end{table}

\subsection{Overall Performance on Validation Set}


Table \ref{tab:overall_performence} shows the results of our method on validation set. Among all the results we post in table, we utilized the same training settings as we described in experiment settsings. As the result we post, different feature combinations can lead to different effects.

\subsection{Model Ensemble}

During the model selection phase, we trained some models with different structures, combinations of features and hyper-parameters, and achieve competitive performence. Vote strategy is employed to improve robustness and perfermence of our final prediction, which is described in table \ref{tab:ensemble}. As the results we post, ensemble different architecture and differnet feature combinations can lead to more benefits on validation set.

\begin{table}[]
\begin{center}
  \caption{The results of each single model and the ensemble of them for the expression prediction task on the validation set.}
  \label{tab:ensemble}
\resizebox{\columnwidth}{!}{
\begin{tabular}{c|c|c}
\hline
Model           & Features                                          & F1                \\ \hline
Transformer     & mae,ires100,wav2vec,ecapatdnn,hubert              & 0.38362           \\
Transformer     & ires100,fau,hubert,wav2vec,ecapatdnn              & 0.35119           \\
Transformer     & ires100,fau,densenet,ecapatdnn,hubert             & 0.36087           \\
Transformer     & mae,ires100,densenet,ecapatdnn,hubert             & 0.39380           \\
LSTM            & densenet,mae,ires100,wav2vec,ecapatdnn,hubert     & 0.40178           \\
LSTM            & mae,ires100,wav2vec,ecapatdnn,hubert              & 0.40832           \\
LSTM            & fau,ires100,ecapatdnn,hubert                      & 0.38928           \\
LSTM            & densenet,mae,ires100,ecapatdnn,hubert             & 0.40889           \\
LSTM            & densenet,mae,ires100,ecapatdnn,hubert             & 0.41410           \\ \hline
Ensemble        &                                                   & \textbf{0.45774}  \\ \hline
\end{tabular}
}
\end{center}
\end{table}

\subsection{Ablation Study}

Table \ref{tab:ablation} shows the abalation study perfermence on validation set, we utilized transformer-based model to compare the benefits of different features and feature combinations. Among all the experimental results, we utilized the same experimental setup in the training phase, except for the feature combinations.

\begin{table}[]
\begin{center}
  \caption{Ablation study of features on the validation set.}
  \label{tab:ablation}
\resizebox{\columnwidth}{!}{
\begin{tabular}{c|c|c}
\hline
Visual                      & Audio                     & F1        \\ \hline
fau,densenet                & ecapatdnn                 & 0.35276   \\ 
fau,densenet                & fbank,ecapatdnn           & 0.33887   \\ 
fau,densenet                & ecapatdnn,hubert          & 0.34391   \\ 
fau,ires100,densenet        & ecapatdnn,hubert          & 0.36266   \\
fau,ires100,densenet        & wav2vec,ecapatdnn         & 0.36944   \\
mae,ires100,densenet        & ecapatdnn,hubert          & 0.39380   \\ \hline
\end{tabular}
}
\end{center}
\end{table}

\section{Conclusion}

In this paper, we propose our framework for the Expression Classification Challenge of the fifth Affective Behavior Analysis in-the-wild (ABAW) Competition. Our approach leverages information from multiple modalities in the spatio-temporal dimension. Various temporal encoders are applied to capture the temporal contextual information in the video. In addition, we design multiple high-quality feature combinations to extract more effective emotional information. Our method achieves a performance of $0.45774$ on the validation set.

\bibliographystyle{unsrt}
\bibliography{main}

\end{document}